OpenStudy (anonymous):

If a die is rolled twice, what is the probability of rolling a 5 and then a 2? 1 36 B) 1 3 Eliminate C) 1 6 D) 2 36

OpenStudy (anonymous):

What is the question.

OpenStudy (texaschic101):

I have no idea what this says

OpenStudy (anonymous):

Me neither

OpenStudy (anonymous):

OHH. thanks

OpenStudy (anonymous):

I meant to type the question too.. Its If a die is rolled twice, what is the probability of rolling a 5 and then a 2?

OpenStudy (mrnood):

OK @kkittykat answer these 3 questions first: are the 2 instances of rolling the dice independent events? (that is for YOU to decide) what is the chance of rolling a 5 with one dice? what is the chance of rolling a 2 with one dice?

OpenStudy (anonymous):

Idk about the independent events and 5 over 6 and 2 over 6 i think

OpenStudy (jameshorton):

So, prob. of rolling a 2: 1/6 You want the second roll to be anything but a 2. So that means you have a choice of 5 sides (1,3,4,5,6) and each of these sides has a 1/6 probability of showing up. Therefore the combined probability of rolling any of these five sides = 5*(1/6) = 5/6 Therefore, the probability of rolling a 2 then anything but a 2: = Prob. of rolling a 2 * Prob. of rolling anything but a 2 = 1/6 * 5/6 = 5/36...Answer

OpenStudy (mrnood):

Well you need to decide about independent events do the results of the first throw influence the results of the second throw? if NOT then the events are independent And the probabilities you gave are wrong there are six possible results on a die - what is the chance of a particular one of those coming up (e.g. a 5?)

OpenStudy (mrnood):

@jameshorton that is not the correct interpretation of the question you answer is not correct

OpenStudy (mrnood):

NO - first - it is against the rules second you do not 'need to get the maths grade' - you need to know HOW to do it if oyu did it th esame way a sjames then you did it wrong like James. answer this first: there are six possible results on a die - what is the chance of a particular one of those coming up (e.g. a 5?)

OpenStudy (mrnood):

going somewhere else for the answer is not 'HELP' if oyu spend your time at school copying other people's work then you are wasting the best opportunity of your life. I am prepared to spend as much time HELPING as you need - but no time giving you th eanswer try to understand the method - then you will have not only the answer to THIS question - but any others like it AND possibly an understanding of chance and odds that may save you a lot of time/money in REAL life. Eductation is YOUR choice - and it is not about grades it's about understanding

OpenStudy (mrnood):

I didn't say 'try again' I gave you 2 clues: first do the results of the first throw influence the results of the second throw? if NOT then the events are independent This is something that you may need to think about - but it is key to working out odds like this. If you throw for instance a 5 with the first throw then you are EQUALLY likely to throw a 5 with th e second throw - i.e. the first does NOT influence the second The chance of throwing a 5 is the same EVERY time oyu throw a die so the question is what is that chance? which is where my second hint comes in: there are six possible results on a die - what is the chance of a particular one of those coming up (e.g. a 5?) i.e what is the chance of any ONE out of SIX events happening?